10 beef animals will start in the feedlot at an average of 350 kg . They will be fed to a slaughter mass of 550kg on average . The present ADG is 0.80kg and this will be improved by 20%. Weaners have already experienced total cost of 16 000 rands when they get introduced into feedlot . Total cost per day in the feedlot are 500 rands . The total income per finished animal sold will be 75 rands per kilogram. Calculate the expected profit from your beef animal according to the improved ADG.
1
Expert's answer
2018-10-01T13:12:09-0400
The number of days to feed is (550 - 350)/(0.8*1.2) = 208.33 days.
The expected profit from your beef animal according to the improved ADG is:
TP = (AR - (AFC + AVC))*Q = (75*550 - (16000/10 + 208.33*500))*10 = -645150, so you will face losses